OpenStudy (anonymous):

Tricky. If x^2+x=3, then x^4+x= ?

OpenStudy (anonymous):

\[x^{2}+x=3, x ^{4}+x= ?\]

OpenStudy (anonymous):

it is a system of equations right? Solve for one and pluge it into the other

OpenStudy (anonymous):

No, MCQs: a) x^2+3 b) 4(3-x) C) 3(x-1) D) 6(2-x)

OpenStudy (anonymous):

still solve for the first and plug in to the second eq

OpenStudy (anonymous):

Can't solve for the first

OpenStudy (anonymous):

I don't know how to solve

OpenStudy (anonymous):

use the quadratic formula

OpenStudy (anonymous):

this is an odd one the options you are given wouldn't work. Maybe it is asking something different but idk what.

OpenStudy (anonymous):

It is 6(2-x)

OpenStudy (anonymous):

Try to see why.

OpenStudy (anonymous):

\[ x^4 + x= x(x^3+1)= x(x+1) ( x^2 - x +1)=\\ (x^2 +x)( (x^2 +x) -x -x +1)= 3( 3 -2x + 1)=3( 4 -2 x)=6(2-x) \]

OpenStudy (anonymous):

Thank you so much Elia, however, I figured this out in the morning today lol.

OpenStudy (anonymous):

Excellent